# Crafting Adversarial Inputs for Large Vision-Language Models Using Black-Box Optimization

### v2

@@ 5 Ablation Study
We conduct ablation studies on two critical parameters of the proposed attack strategy: (1) Analyzing the impact of finite difference noise magnitude $\boldsymbol{\Delta}$ to $\mathrm{e}^{-5}$. (2) Evaluating the effectiveness of Rademacher-distributed noise for gradient estimation.
@@
Table 7: Toxicity metrics (%) across different numbers of iterations with a fixed noise step size (e.g., K = 1,000 iterations).
(1) Impact of finite difference noise magnitude. For noise values, this hyperparameter determines the sensitivity of SPSA, directly influencing gradient estimation resolution. Smaller noise values enable finer-grained gradient estimation but introduce a trade-off: excessively small step sizes can lead to numerical instability and potentially weaken attack effectiveness due to vanishing differential signals. Table 7 shows how toxicity scores from the Perspective API and the Detoxify classifier vary across six toxicity attributes under ZO-SPSA with smaller Gaussian noise. As iteration count increases, both evaluators show lower toxicity rates for six attributes. This trend becomes less prominent when using a smaller noise magnitude, suggesting that smaller noise values undermine attack effectiveness. These observations emphasize the importance of properly adjusting ZO-SPSA noise parameters for optimal adversarial effectiveness.
@@
Table 8: Toxicity metrics (%) under varying numbers of iterations (e.g., K = 1,000 iterations) with Rademacher-distributed perturbations.
(2) Rademacher-distributed noise for gradient estimation accuracy. This approach uses a randomized central finite difference scheme, where the binary symmetric noise variable is either +1+1 or 1-1. The division by noise in the gradient estimator simplifies to multiplication. The formulation is shown in Eq. 3:
@@
where $\boldsymbol{\Delta}$ represents a random noise vector with each component $\Delta_{i}$ following a Rademacher distribution (taking values of +1 or -1 with equal probability) to provide random noise directions for gradient estimation.
Table 8 shows the ZO-SPSA attack using Rademacher-distributed noise under different iterations. Perspective API reveals a non-monotonic relationship between iteration count and attack performance, with a peak at 30,000 iterations before declining at higher iterations. Specifically, the toxicity rate for the “Any” attribute drops from 42.2% at 5,000 iterations to 35.2% at 40,000 iterations, with a sharp peak of 49.7% observed at 30,000 iterations. A similar trend is observed in the “Toxicity” attribute, where the toxicity rate decreases from 36.3% to 29.9%, with a notable spike to 43.5% at 30,000 iterations. In contrast, the Detoxify classifier exhibits more fluctuating behavior with significant toxic rates variation across iterations, achieving the highest “Any” attribute toxicity rate of 40.3% at 30,000 iterations. These findings indicate that ZO-SPSA with Rademacher-distributed noise shows optimal effectiveness at specific iteration counts and exhibits inconsistent performance across different iteration settings.
@@ 6 Conclusion
Our work has the following limitations. First, the method requires extensive forward propagation of the target model, which leads to computational inefficiency. Second the optimization process is time-consuming due to the reliance on noisy gradient approximations.

### v3

@@ 5 Ablation Study
We conduct ablation studies on two critical parameters of the proposed attack strategy: (1) Analyzing the impact of finite difference perturbation magnitude $\boldsymbol{\Delta}$ to $\mathrm{e}^{-5}$. (2) Evaluating the effectiveness of Rademacher-distributed perturbation for gradient estimation.
@@
Table 7: Toxicity metrics (%) reported by Perspective API and Detoxify under a fixed perturbation scale of 1e-5 and varying iteration steps (e.g., K = 1,000 iterations).
(1) Impact of finite difference perturbation magnitude. This hyperparameter determines the sensitivity of SPSA, directly influencing gradient estimation resolution. Smaller perturbation values enable finer-grained gradient estimation but introduce a trade-off: excessively small step sizes can lead to numerical instability and potentially weaken attack effectiveness due to vanishing differential signals. Table 7 presents a sensitivity analysis of updating SPSA iteration counts (5,000 to 40,000) under fixed perturbation step size ($\delta=1\times 10^{-5}$) to assess how Perspective API and Detoxify classifier vary across six toxicity attributes. As iteration count increases, both evaluators show lower toxicity rates for six attributes. This trend becomes less prominent when using a smaller perturbation magnitude, suggesting that smaller perturbation values undermine attack effectiveness. These observations emphasize the importance of properly adjusting ZO-SPSA perturbation parameters for optimal adversarial effectiveness.
@@
Table 8: Toxicity metrics (%) reported by Perspective API and Detoxify under varying numbers of iterations (e.g., K = 1,000 iterations) with Rademacher-distributed perturbations.
(2) Rademacher-distributed perturbation for gradient estimation accuracy. This approach uses a randomized central finite difference scheme, where the binary symmetric noise variable is either +1+1 or 1-1. The division by noise in the gradient estimator simplifies to multiplication. The formulation is shown in Eq. 3:
@@
where $\boldsymbol{\Delta}$ represents a random perturbation vector with each component $\Delta_{i}$ following a Rademacher distribution (taking values of +1 or -1 with equal probability) to provide random noise directions for gradient estimation.
Table 8 shows the ZO-SPSA attack using Rademacher-distributed perturbation under different iterations. Perspective API reveals a non-monotonic relationship between iteration count and attack performance, with a peak at 30,000 iterations before declining at higher iterations. Specifically, the toxicity rate for the “Any” attribute drops from 42.2% at 5,000 iterations to 35.2% at 40,000 iterations, with a sharp peak of 49.7% observed at 30,000 iterations. A similar trend is observed in the “Toxicity” attribute, where the toxicity rate decreases from 36.3% to 29.9%, with a notable spike to 43.5% at 30,000 iterations. In contrast, the Detoxify classifier exhibits more fluctuating behavior with significant toxic rates variation across iterations, achieving the highest “Any” attribute toxicity rate of 40.3% at 30,000 iterations. These findings indicate that ZO-SPSA with Rademacher-distributed perturbation shows optimal effectiveness at specific iteration counts and exhibits inconsistent performance across different iteration settings.
@@ 6 Conclusion
Our work has the following limitations. First, the method requires requires extensive forward propagation of the target model, which leads to computational inefficiency. Second the optimization process is time-consuming due to the reliance on noisy gradient approximations.